A man borrows Rs. 360 if he pays it back in 12 monthly installments of Rs. 31.50, what is his interest rate?
Found 2 solutions by fractalier, Shin123:
Answer by fractalier(6550)

Well 12 payments of 31.50 adds up to Rs 378.
That means he paid 378 - 360 = Rs 18 interest.
18/360 = 5% interest

Answer by Shin123(626)

31.50*12=378 378-360=18 18/360=1/20=0.05=5% 5% interest